# Natal, Brazil

### v1

@@ Abstract
One of the crucial activities in software development is monitoring. It plays a vital role in verifying the proper implementation of processes, the identification of errors, and the discovery of opportunities for improvement. Continuous Integration (CI) encompasses a set of widely adopted practices that enhance software development. However, there are indications that developers may not adequately monitor CI practices. Hence, this paper explores developers’ perceptions regarding the monitoring CI practices. To achieve this, we first perform a Document Analysis to assess developers’ expressed need for practice monitoring in pull requests comments generated by developers during the development process. After that, we conduct a survey among developers from 121 open-source projects to understand perception of the significance of monitoring seven CI practices in their projects. Finally, we triangulate the emergent themes from our survey by performing a second Document Analysis to understand the extent of monitoring features supported by existing CI services. Our key findings indicate that: 1) the most frequently mentioned CI practice during the development process is “Test Coverage” (> 80%), while “Build Health” and “Time to Fix a Broken Build” present notable opportunities for monitoring CI practices; 2) developers do not adequately monitor all CI practices and express interest in monitoring additional practices; and 3) the most popular CI services currently offer limited native support for monitoring CI practices, requiring the use of third-party tools. Our results lead us to conclude that monitoring CI practices is often overlooked by both CI services and developers. Using third-party tools in conjunction with CI services is challenging, they monitor some redundant practices and still falls short of fully supporting CI practices monitoring. Therefore, CI services should implement CI practices monitoring, which would facilitate and encourage developers to monitor them.
@@ 1 Introduction
Continuous Integration (CI) is a set of software development practices where change sets that teams produce are built and tested as they are submitted for inclusion in a repository of record. In many settings, CI is invoked several times per day. Each invocation triggers an automated build, which can ideally detect errors as early as possible Duvall2007 . Over the past two decades, there has been a growing acknowledgement of the benefits derived from increasing the frequency of integration and release activities. The term “Continuous ” (continuous star) was introduced by Fitzgerald et al. fitzgerald2014continuous to describe this collection of continuous activities.
@@
Although CI has become a popular topic in research and industry StakeholderPerceptions2015 ; ReplicationCIPainPoints2019 ; TradeOffsCI2017 ; ImpactCICodeReviews2020 ; hilton2016usage ; benefitsCIIndustry2013 ; elazhary2021uncovering ; VasilescuQualityProdOutcomes2015 ; JoaoHelis2018 ; santos2022investigating ; ghaleb2019empirical ; coverageImpac2017 , it is not always the case that development teams fully embrace the spirit of CI, and as such, tend not to enjoy its full benefits CITheater2019 . Development teams may perform this so-called “CI Theater” due to the costs incurred by and challenges encountered when adopting CI. Indeed, Hilton et al. hilton2016usage and Elazhary et al. elazhary2021uncovering asked developers why they did not (fully) adopt CI, observing recurrent themes, such as, tradeoff between test effectiveness and build speed ghaleb2019empirical .
@@
Despite the central role that CI plays in software development, the extent to which stakeholders are aware of the maturity of their CI workflows remains unclear. We believe that such an awareness would require active monitoring of the CI practices. For example, a company studied by Elazhary et al. elazhary2021uncovering admittedly runs tests inefficiently due project constraints on the available test infrastructure. In turn, such choices prolong build durations, which can hinder organizational agility chen2020buildfast . However, what if the monitoring of build durations, could help said company make the business case to invest in additional test infrastructure? Without awareness of the current CI workflow, such investment decisions tend to be made intuitively, and as such, are prone to limitations and biases. Indeed, the recent debut of Metrics GitHub Actions service  [^1] https://github.blog/2023-07-19-metrics-for-issues-pull-requests-and-discussions/ suggests that monitoring metrics related to CI is a feature that CI providers believe is practical and relevant.
@@
As such, our research focuses on a set of questions regarding the monitoring of CI practices, such as: do development teams monitor their CI practices? Do they perceive the monitoring of CI practices as important? Which monitoring capabilities do state-of-the-art CI services provide? Continuous monitoring facilitates the timely identification of quality-of-service issues, including performance degradation, as well as the fulfillment of Service Level Agreements (SLAs) fitzgerald2014continuous . Typically, monitoring focuses on application performance (e.g, CPU, memory, network, and disk) and infrastructure (e.g, physical servers, virtual machines, databases, network infrastructure), nonetheless our work focuses on monitoring the software integration process, which is often overlooked.
@@
To better understand the importance and prevalence of monitoring in the context of CI practices, we conduct a series of qualitative analyses. We start by investigating whether developers express the need for monitoring CI practices in their project repositories. To do so, in RQ1, we perform a Document Analysis bowen2009documentanalysis on 4,892 Pull-Requests(PR) comments to identify the situations where developers have expressed the need for monitoring CI practices.
@@
  • RQ1: Do developers express the need to monitor CI practices in PRs? Our findings show that the main needs expressed by developers are related to monitoring “build health” and “time to fix a broken build” since they often encounter build failures in PR comments. We also discover that “Coverage” is the most monitored practice (83% of the monitoring opportunities found in PR comments, suggested the monitoring of Coverage), and the CodeCov tool is the most commonly used tool for monitoring this practice (86% of coverage monitoring in PR comments used this tool).
@@
  • RQ3: Which monitoring features are currently supported by existing CI tools? CI services currently offer limited built-in support for monitoring CI. Only the Gitlab documentation suggests an implementation of a dashboard for monitoring practices related to continuous activities. Build metadata tracking is the most commonly supported monitoring practice. There are various complementary tools that can enhance this monitoring. However, they often come with additional costs and complexity, with a predominant focus on fundamental information about the performance of the CI pipeline.
Our findings highlight that “Coverage” was the most frequently monitored practice $258306\frac${258}{306}divide start_ARG 258 end_ARG start_ARG 306 end_ARG ( or 80%), with CodeCov being the main tool used in most projects $206258\frac${206}{258}divide start_ARG 206 end_ARG start_ARG 258 end_ARG ( or 79%), a result reinforced by the participants of our survey. The primary opportunities lie in monitoring “build health” and “time to fix a broken build”. A significant majority of participants in our survey $2028\frac${20}{28}divide start_ARG 20 end_ARG start_ARG 28 end_ARG ( or 71%) demonstrated interest in monitoring CI practices. Furthermore, we observed that CI services do not provide built-in support to monitor CI practices and the adoption of third-party tools for monitoring CI practices still presents challenges that need to be overcome.
@@ 2.1 CI Adoption and Its Impacts
There is a considerable body of literature on the topic of CI adoption, exploring its benefits and costs. However, most of these studies have concentrated on evaluating CI adoption in its entirety (i.e., evaluating effects prior and after adopting CI). Many of these studies lack clear criteria for defining the use of CI  EliezioRevisaoSistematica2022 . Among the studies lacking precise criteria, the most commonly applied criterion, based on interviews, surveys, or other data sources (e.g., build logs), was the adoption of a CI service.
@@ 2.2 CI Practices and Its Impacts
Duvall et al. Duvall2007 recommended seven best practices for individuals and teams running CI on a project: i) Commit Code Frequently, ii) Do not Commit Broken Code, iii) Fix Broken Builds Immediately, iv) Write Automated Developer Tests, v) All Tests and Inspections must Pass, vi) Run Private Builds, and vii) Avoid Getting Broken Code. While recommendations are plentiful, there is still no consensus on the set of CI practices that a project should follow, and what specific values teams should strive for (e.g., number of commits per day or hour). Indeed, the ideal extent to which these practices are employed may vary from project to project, depending on their domain, the quantity of features being produced/maintained, and/or the size of its user base. This variation in CI adoption makes monitoring even more challenging. In this paper, we analyze the following measures that align with Duvall et al.’s Duvall2007 seven best practices for CI: (i) Commit Per Weekday, (ii) Coverage, (iii) Build Duration, (iv) Build Activity/Build-frequency, (v) Build Health/Build Quality, (vi) Time to Fix a Broken Build, and (vii) Comments Per Change/Comments Per PR.
Santos et al. santos2022investigating collected data on CI practices from 90 relevant open-source projects over a 2-year period and performed a quantitative analysis to investigate whether these practices are related to higher project productivity or quality. Furthermore, Elaszhary et al. elazhary2021uncovering investigated the extent to which three software organizations implement ten Continuous Integration (CI) practices defined by Fowler fowler2006continuous . They explored the benefits these practices bring and the challenges encountered during their implementation. This inquiry was conducted through a multiple-case study with mixed methods, focusing on three small to medium-sized software-as-a-service organizations.
@@ 3 Methodology
In this section, we present our research questions. For each questions, we explain our motivation and methodological approach.
@@
Motivation: The starting point of our research is to understand whether developers express their concern regarding monitoring the adherence level to the CI practices in their projects. Understanding developers’ concerns can help the developers to identify areas for improvement and ensure that best practices are being followed. Moreover, this can lead to further research initiatives aimed at developing better tools, methodologies, and best practices for CI monitoring in software development projects. We analyze the developer discussion reflected on the comments of pull requests (PRs). We have two main goals with this analysis:
@@
Approach: To collect a representative sample of PRs comments that provide insight into how developers discuss CI monitoring, we first need to select a representative sample of non-trivial open-source projects. Following an approach similar to other studies VasilescuQualityProdOutcomes2015 ; Vassallo01 ; JoaoHelis2018 ; santos2022investigating , we outlined the selection process steps in Figure 1. In the following paragraphs, we will detail the steps taken to select the projects for our study.
@@
In Step 1 of our study, we search for the most starred repositories with more than 100 stars on GitHub across a set of 15 popular programming languages: Java, JavaScript, C#, Python, PHP, TypeScript, C, Go, C++, Kotlin, Ruby, Rust, Swift, Scala, and Objective-C. We exclude repositories that were forks and that were smaller than 10MB. Our search yielded a pool of 41,469 repositories. We filter out repositories with no more than 100 pull requests (PRs) and 100 issues, resulting in a reduced set of 11,014 projects. The threshold value of 100 was chosen based on previous studies VasilescuQualityProdOutcomes2015 ; JoaoHelis2018 ; santos2022investigating . Our threshold of 10MB for repository size was also based on prior research MarcosOliverira2017 ; DEOLIVEIRA2019110420 ; ConvolutionalAttentionNetwork2017 ; santos2022investigating which has demonstrated that repositories smaller than 10MB are unlikely to represent meaningful software projects.
In Step 2, we check whether repositories used the TravisCI or GitHub Actions (GH-Actions) services. We focus on these two CI services because they are the most widely used CI services in open-source projects, with TravisCI having a usage rate of 44% and GH-Actions a usage rate of 39% GolzadehCIServer2022 . However, when we checked the build information of these CI services, only three repositories using TravisCI had sufficient build information for our analyses. Aside from that, given the significantly decrease in the adoption rate of TravisCI DecanGitHubAction2022 , we decide to focus solely on repositories using GH-Actions for our analyses, i.e., we discarded repositories using TravisCI. We ended up with 7,658 GH-Actions repositories at the end of this step.
Unlike other CI services, GH-Actions can define several workflows that can be responsible for executing tasks unrelated to CI. The documentation of GH-Actions states the following: “A workflow is an automated process composed of a series of jobs that gets executed when it’s triggered by an event. Workflows are defined in YAML files and are stored in a .github/workflows directory at the root of the repository. A repository can also have multiple workflows” GHActionGetStarted2013 . It was, therefore, necessary to identify which workflows were related to CI in Step 3 and focus our analysis only on these workflows. To do so, we considered the following criteria: 222our scripts to identify CI-related workflows is publicly available on https://github.com/jadsonjs/ci-workflows-trader.
@@
Using the criteria above, we find an average of 2.89 CI workflows per project, with a maximum of 65, a minimum of 1, and a standard deviation of 3.1.
@@
In Step 4, we check if projects have been using CI for at least 2 years (last build - first build on GH-Action >= 24 months) and if the last build happened in less than 6 months (last build <= 6 months). GH-Action does not have the “build” concept as, for example, TravisCI does. We, therefore, consider a build to be a Runner of a CI workflow, which was considered in Step 3. In fact, GH-Actions define runners as: “processes on a server that run the workflow when it’s triggered”GHActionGetStarted2013 . We refer to Runners as builds for the sake of simplicity. We obtain a total of 2,488 projects with CI-related workflows.
@@
In Step 5, we collect information related to builds, issues, pull requests, commits, pull request comments, and code coverage within the period of the last two years of CI. Next, we filter for projects that have at least 100 builds, 100 issues, 100 pull requests, 100 commits, and 100 coverage entries. As for coverage, we focus on the Coveralls [^3] https://docs.coveralls.io/api-introduction and CodeCov [^4] https://docs.codecov.com/reference/authorization APIs, as they are two of the most popular coverage tools kavaler2019tool . In total, we collected 871,104 pull request comments, 405,909 pull requests, 592,812 issues, 766,214 commits and 2,595,954 builds from 474 projects.
@@
In Step Step 6, we filter for projects that have build information for each of the last 24 months (last build - 24 months), resulting in a final set of 121 projects. These projects together have a total sum of 73,004 PRs that were used to answer these research questions. We created a representative sample SampleSize2021 of 382 randomly selected PRs (95% confidence level and 5% confidence interval) that contain at least one comment. Subsequently, we collected all the comments and reviews of these PRs, resulting in a total of 4,892 comments. We performed a Thematic Analysis maguire2017doingThematicAnalysis to identify needs and evidence of current monitoring practices. While these comments are not exclusively about CI practices, they contain genuine discussions between developers, which may represent concerns about CI practices if they are pertinent to the project. The extracted codes and themes were initially assigned by the first author. To reduce the probability of errors inherent in this type of subjective analysis, the second author independently applied the coding guide to label all examples. Finally, we calculated the agreement score between the two sets of labels using the Cohen Kappa statistical coefficient cohen1968weighted , which measures the agreement between two “judges”. We obtained a score of 0.68 (94.94% of agreement), indicating “Substantial Agreement”.
@@
Motivation: In light of the observed level of concern among developers regarding monitoring CI practices in RQ1, we delve deeper into this inquiry in RQ2 to directly solicit developers’ opinions on the importance of monitoring CI practices conducting a survey. We aim to address questions such as: Do developers prioritize monitoring their CI practices? Do they adhere to specific monitoring processes? Through RQ2, we seek to understand developers’ perceptions of CI practice monitoring and check whether they recognize its importance. Additionally, encourage discussion, knowledge sharing, and community building around best CI practices.
@@
In Step 8 of Figure 1, to identify bug-related issues, we first collected labels associated with projects issues. This approach is similar to Vasilescu et al. VasilescuQualityProdOutcomes2015 , where we manually check our projects to understand how they use labels to indicate the presence of bugs. We download all the labels from the 121 projects and checked their description to understand whether the label represented a bug. The resulting list of labels representing bug-related issues is: “defect”, “error”, “bug”, “issue”, “mistake”, “incorrect”, “fault”, “flaw”, “crash”, “bugfix” and “regression.” Similar to the work of Vasilescu et al., VasilescuQualityProdOutcomes2015 , we lowercased the labels and apply the Porter stemming algorithm  [^5] https://tartarus.org/martin/PorterStemmer/ to compare the bug-related labels. Finally, bug-related issues were analyzed to determine correlations between CI practices and the number of bugs in the projects. This data was used in an exploratory survey question to ascertain whether developers observed any correlations between CI practices and the incidence of bugs in the project.

### v2

@@ Abstract
One of the crucial activities in software development is monitoring. It plays a vital role in verifying the proper implementation of processes, the identification of errors, and the discovery of opportunities for improvement. Continuous Integration (CI) encompasses a set of widely adopted practices that enhance software development. However, there are indications that developers may not adequately monitor CI practices. Hence, this paper explores developers’ perceptions regarding monitoring CI practices. To achieve this, we first defined metrics associated with CI practices. Then, we conducted a series of analyses to evaluate how these metrics are perceived in projects or by their developers. As a starting point, we perform a Document Analysis to assess developers’ expressed need for practice monitoring in pull request comments generated by developers during the development process. After that, we conduct a survey of 28 developers from 121 open-source projects to understand the perception of the significance of monitoring seven CI practices in their projects. Finally, we triangulate the emergent themes from our survey by performing a second Document Analysis to understand the extent of monitoring features supported by existing CI services. Our key findings indicate that: 1) the most frequently mentioned CI practice during the development process is “Write automated developer tests”, which is associated with the metric “Coverage” (> 80%), while “Don’t commit broken code” and “Fix Broken Builds Immediately’ present notable opportunities for monitoring CI practices; 2) developers do not adequately monitor all CI practices and express interest in monitoring additional practices; and 3) the most popular CI services currently offer limited native support for monitoring CI practices, requiring the use of third-party tools. Our results lead us to conclude that monitoring CI practices is often overlooked by both CI services and developers. Using third-party tools in conjunction with CI services is challenging, they monitor some redundant practices, and they still fall short of fully supporting CI practices monitoring. Therefore, CI services should implement CI practice monitoring, which would facilitate and encourage developers to monitor them.
@@ 1 Introduction
Continuous Integration (CI) is a set of software development practices where change sets that teams produce are built and tested as they are submitted for inclusion in a repository of records. In many settings, CI is invoked several times per day. Each invocation triggers an automated build, which can ideally detect errors as early as possible Duvall2007 . Over the past two decades, there has been a growing acknowledgement of the benefits derived from increasing the frequency of integration and release activities. The term “Continuous ” (continuous star) was introduced by Fitzgerald et al. fitzgerald2014continuous to describe this collection of continuous activities.
@@
Although CI has become a popular topic in research and industry StakeholderPerceptions2015 ; ReplicationCIPainPoints2019 ; TradeOffsCI2017 ; ImpactCICodeReviews2020 ; hilton2016usage ; benefitsCIIndustry2013 ; elazhary2021uncovering ; VasilescuQualityProdOutcomes2015 ; JoaoHelis2018 ; santos2022investigating ; ghaleb2019empirical ; coverageImpac2017 , it is not always the case that development teams fully embrace the spirit of CI and, as such, tend not to enjoy its full benefits. Development teams may perform this so-called “CI TheaterCITheater2019 where they have the illusion of practicing CI without fully adopting all its practices, often due to the costs and challenges associated with implementing CI. This lack of complete adoption of CI practices may lead to unhealthy practices, as discussed by Felidré et al. CITheater2019 . Indeed, Hilton et al. hilton2016usage and Elazhary et al. elazhary2021uncovering asked developers why they did not (fully) adopt CI, observing recurring themes, such as the trade-off between test effectiveness and build speed ghaleb2019empirical .
@@
Despite the central role that CI plays in software development, it remains unclear to what extent stakeholders are aware of the maturity of their CI workflows. We believe that such an awareness would require active monitoring of CI practices. For example, a company studied by Elazhary et al. elazhary2021uncovering admittedly runs tests inefficiently due to project constraints on the available test infrastructure. In turn, such choices prolong build durations, which can hinder organizational agility chen2020buildfast . However, what if the monitoring of build durations could help the said company make the business case to invest in additional test infrastructure? Without awareness of the current CI workflow, such investment decisions tend to be made intuitively and, as such, are prone to limitations and biases. Indeed, the recent debut of Metrics GitHub Actions service  [^1] https://github.blog/2023-07-19-metrics-for-issues-pull-requests-and-discussions/ suggests that monitoring metrics related to CI is a feature that CI providers believe is practical and relevant.
@@
As such, our research focuses on a set of questions regarding the monitoring of CI practices, such as: do development teams monitor their CI practices? Do they perceive the monitoring of CI practices as important? Which monitoring capabilities do state-of-the-art CI services provide? Continuous monitoring facilitates the timely identification of quality-of-service issues, including performance degradation, and the fulfillment of Service Level Agreements (SLAs) fitzgerald2014continuous . Typically, monitoring focuses on application performance (e.g, CPU, memory, network, and disk) and infrastructure (e.g, physical servers, virtual machines, databases, network infrastructure), nonetheless, our work focuses on monitoring the software integration process, which is often overlooked.
@@
To better understand the importance and prevalence of monitoring in the context of CI practices, we conduct a series of qualitative analyses. We start by investigating whether developers express the need for monitoring CI practices in their project repositories. To do so, in RQ1, we perform a Document Analysis bowen2009documentanalysis on 4,892 Pull-Request (PR) comments to identify the situations where developers have expressed the need for monitoring CI practices.
@@
  • RQ1: Do developers express the need to monitor CI practices in PRs? Our findings show that the main needs expressed by developers are related to monitoring “Don’t commit broken code” and “Fix Broken Builds Immediately” since they often encounter build failures in PR comments. We also discover that “Write Automated Developer Tests” is the most monitored practice (83% of the monitoring opportunities found in PR comments, suggested the monitoring of Coverage metric), and the CodeCov tool is the most commonly used tool for monitoring this practice (86% of coverage monitoring in PR comments used this tool).
@@
  • RQ3: Which monitoring features are currently supported by existing CI tools? CI services currently offer limited built-in support for monitoring CI. Only the Gitlab documentation suggests an implementation of a dashboard for monitoring metrics related to continuous activities. Build metadata tracking are the most commonly supported monitoring metrics. There are various complementary tools that can enhance this monitoring. However, they often come with additional costs and complexity, with a predominant focus on fundamental information about the performance of the CI pipeline.
Our findings highlight that “Coverage” (that is associated with “Write Automated Developer Tests” practice) was the most frequently monitored metric $258306\frac${258}{306}divide start_ARG 258 end_ARG start_ARG 306 end_ARG ( or 80%), with CodeCov being the main tool used in most projects $206258\frac${206}{258}divide start_ARG 206 end_ARG start_ARG 258 end_ARG ( or 79%), a result reinforced by the participants of our survey. The primary opportunities lie in monitoring “Don’t commit broken code” and “Fix Broken Builds Immediately”. A significant majority of participants in our survey $2028\frac${20}{28}divide start_ARG 20 end_ARG start_ARG 28 end_ARG ( or 71%) demonstrated interest in monitoring CI practices. Furthermore, we observed that CI services do not provide built-in support to monitor CI practices and the adoption of third-party tools for monitoring CI practices still presents challenges that need to be overcome.
@@ 2.1 CI Adoption and Its Impacts
There is a considerable body of literature on the topic of CI adoption, exploring its benefits and costs. However, most of these studies have concentrated on evaluating CI adoption in its entirety (i.e., evaluating effects prior to and after adopting CI). Many of these studies lack clear criteria for defining the use of CI  EliezioRevisaoSistematica2022 . Among the studies lacking precise criteria, the most commonly applied criterion, based on interviews, surveys, or other data sources (e.g., build logs), was the adoption of a CI service.
@@ 2.2 CI Practices and Its Impacts
Duvall et al. Duvall2007 recommended seven best practices for individuals and teams running CI on a project: i) Commit Code Frequently, ii) Do not Commit Broken Code, iii) Fix Broken Builds Immediately, iv) Write Automated Developer Tests, v) All Tests and Inspections must Pass, vi) Run Private Builds, and vii) Avoid Getting Broken Code. While recommendations are plentiful, there is still no consensus on the set of CI practices that a project should follow, and what specific values teams should strive for (e.g., number of commits per day or hour). Indeed, the ideal extent to which these practices are employed may vary from project to project, depending on their domain, the quantity of features being produced/maintained, and/or the size of its user base. This variation in CI adoption makes monitoring even more challenging. In this paper, we analyze the following measures that align with Duvall et al.’s Duvall2007 seven best practices for CI: (i) Commit Per Weekday, (ii) Coverage, (iii) Build Duration, (iv) Build Activity/Build-frequency, (v) Build Health, (vi) Time to Fix a Broken Build, and (vii) Comments Per Change/Comments Per PR.
Santos et al. santos2022investigating collected data on CI practices from 90 relevant open-source projects over a 2-year period and performed a quantitative analysis to investigate whether these practices are related to higher project productivity or quality. Furthermore, Elaszhary et al. elazhary2021uncovering investigated the extent to which three software organizations implement ten Continuous Integration (CI) practices defined by Fowler fowler2006continuous . They explored the benefits these practices bring and the challenges encountered during their implementation. This inquiry was conducted through a multiple-case study with mixed methods, focusing on three small to medium-sized software-as-a-service organizations. Their study reveal which of the ten CI practices the three companies adopted, how and why they adopted them, the benefits they perceived, and the challenges they encountered. Moreover, their findings emphasizes the importance of the context in which CI is applied. Many aspects of CI vary depending on how it is implemented. While organizations broadly adopt CI practices, they adapt them based on perceived benefits, project context, and the CI tools they use. The study maps out the benefits and challenges each company faces when adopting CI practices, focusing on areas where companies differ significantly and where these differences reveal important considerations. Among the key findings, the paper highlights: (i) addressing merge conflicts has a higher priority than the unclear benefits of using a single source repository, (ii) the perceived effort required for integration testing outweighed its perceived benefits, (iii) data-centric projects tend to have longer builds, as comprehensive testing takes priority over shorter build times, and (iv) for some organizations, enforcing deployment privileges to enhance security was prioritized over fostering a sense of ownership in the deployment process. The authors emphasize the need for researchers to consider these variations when studying projects that claim to use CI. Their work also underscores the importance of measuring CI practices to better understand the context in which they are adopted.
@@ 3 Methodology
In this section, we present our research questions. For each question, we explain our motivation and methodological approach.
@@
Motivation: The starting point of our research is to understand whether developers express their concern regarding monitoring the adherence level to the CI practices in their projects. Understanding developers’ concerns can help the developers to identify areas for improvement and ensure that best practices are being followed. Moreover, this can lead to further research initiatives aimed at developing better tools, methodologies, and best practices for CI monitoring in software development projects. To identify CI practices, we examine the CI metrics associated with them in developer discussions reflected in pull request (PR) comments. We have two main goals with this analysis:
@@
We collected data between August and October 2022.
Approach: To collect a representative sample of PR comments that provide insight into how developers discuss CI monitoring, we first need to select a representative sample of non-trivial open-source projects. Following an approach similar to other studies VasilescuQualityProdOutcomes2015 ; Vassallo01 ; JoaoHelis2018 ; santos2022investigating , we outlined the selection process steps in Figure 1. In the following paragraphs, we will detail the steps taken to select the projects for our study.
@@
In Step 1 of our study, we search for the most starred repositories with more than 100 stars on GitHub across a set of 15 popular programming languages: Java, JavaScript, C#, Python, PHP, TypeScript, C, Go, C++, Kotlin, Ruby, Rust, Swift, Scala, and Objective-C. This can be achieved using the GitHub Search REST API [^2] https://docs.github.com/en/rest/search/search, which allows retrieving projects in a specific language, ordered by the number of stars. For this study, we used the Snooper project [^3] https://github.com/jadsonjs/snooper, which provides several pre-imple-mented queries and abstracts calls to the GitHub REST API. We exclude repositories that were forks and that were smaller than 10MB. Our search yielded a pool of 41,469 repositories. We filter out repositories with no more than 100 pull requests (PRs) and 100 issues, resulting in a reduced set of 11,014 projects. We applied these filters to ensure the selection of relevant projects with sufficient activity and data for our study, thus we focused on projects that provided enough information to analyze CI practices meaningfully. The threshold value of 100 was chosen based on previous studies VasilescuQualityProdOutcomes2015 ; JoaoHelis2018 ; santos2022investigating . Our threshold of 10MB for repository size was also based on prior research MarcosOliverira2017 ; DEOLIVEIRA2019110420 ; ConvolutionalAttentionNetwork2017 ; santos2022investigating which has demonstrated that repositories smaller than 10MB are unlikely to represent meaningful software projects.
In Step 2, we check whether repositories used the GitHub Actions
(GHActions) service. We focus on this CI service because it is the most widely used CI service in open-source projects, with a usage rate of 51.7% GolzadehCIServer2022 . Furthermore, GHActions has become the most dominant CI/CD service only 18 months after its introduction, while other CI/CD services, especially Travis, have experienced decreasing adoption and increasing discontinuation rates DecanGitHubAction2022 .
Unlike other CI services, GHActions can define several workflows that can be responsible for executing tasks unrelated to CI. The documentation of GHActions states the following: “A workflow is an automated process composed of a series of jobs that gets executed when it’s triggered by an event. Workflows are defined in YAML files and are stored in a .github/workflows directory at the root of the repository. A repository can also have multiple workflows” GHActionGetStarted2013 . It was, therefore, necessary to identify which workflows were related to CI in Step 3 and focus our analysis only on these workflows. To do so, we considered the following criteria: 444our scripts to identify CI-related workflows is publicly available on https://github.com/jadsonjs/ci-workflows-trader.
@@
Using the criteria above, we identified a total of 19,142 workflows across 6,607 projects. The number of CI workflows per project has a range from 1 to 65. The first quartile (Q1) is 1, the third quartile (Q3) is 3, and the interquartile range (IQR) or median is 2. Notably, the distribution of workflows per project is skewed, with a small number of projects having a disproportionately high number of workflows.
@@
In Step 4, we check if projects have been using CI for at least 2 years (last build - first build on GHActions >= 24 months) and if the last build happened in less than 6 months (last build <= 6 months). GHActions does not have the “build” concept as, for example, TravisCI does. We, therefore, consider a build to be a Runner of a CI workflow, which was considered in Step 3. In fact, GHActions define runners as: “processes on a server that run the workflow when it’s triggered”GHActionGetStarted2013 . We refer to Runners as builds for the sake of simplicity. We obtain a total of 2,488 projects with CI-related workflows.
@@
In Step 5, we collect information related to builds, issues, pull requests, commits, pull request comments, and code coverage within the period of the last two years of CI. Next, we filter for projects that have at least 100 builds, 100 issues, 100 pull requests, 100 commits, and 100 coverage entries. As for coverage, we focus on the Coveralls [^5] https://docs.coveralls.io/api-introduction and CodeCov [^6] https://docs.codecov.com/reference/authorization APIs, as they are two of the most popular coverage tools kavaler2019tool . In total, we collected 871,104 pull request comments, 405,909 pull requests, 592,812 issues, 766,214 commits and 2,595,954 builds from 474 projects. This information was used to calculate CI metrics and identify the use of the evaluated CI practices in the projects.
In Step 6, we filter for projects that have build information for each of the last 24 months (last build - 24 months), to select active projects and prevent projects with CI Theater CITheater2019 . This criterion ensured that the selected projects likely implemented CI practices rather than merely using a CI service (e.g., GitHub Actions) EliezioRevisaoSistematica2022 . Furthermore, since our analysis relied on build information, it was essential for these projects to include such data. This filter resulted in a final set of 121 projects. These projects together have a total sum of 73,004 PRs that were used to answer these research questions. We created a representative sample SampleSize2021 of 382 randomly selected PRs (95% confidence level and 5% confidence interval) that contain at least one comment and were closed within the 24-month period during which data was collected for the study. It is important to note that the selected PRs were not stratified across the 121 projects, meaning they were chosen randomly without division or bias towards any specific project. Subsequently, we collected all the comments and reviews of these PRs, resulting in a total of 4,892 comments. PR comments generated by automated tools, such as bots were indeed included in our sample because we believe these comments may contain useful information generated by CI tools, and for this reason, they were not disregarded. Our random selection process obtained PR comments from 83 of the 121 final projects in our database. Figure 2 presents some contextual factors of the projects in our database.
Figure 2: Contextual Factors of Projects in Our Database
We performed a Document Analysis bowen2009documentanalysis following by Thematic Analysis maguire2017doingThematicAnalysis to identify needs and evidence of current monitoring practices (identifying metrics associated with these CI practices) in PR comments. We believe that Document Analysis combined with Thematic Analysis is the most appropriate choice for analyzing the data due to the large volume and richness of our dataset, which includes approximately 8GB of data including 871,104 PR comments. This extensive collection provides a unique opportunity to extract valuable insights into the developer processes across multiple projects. By systematically examining this large dataset, we can identify patterns, recurring themes, and behaviors related to how CI practices are monitored in these projects. While these comments are not exclusively about CI practices, they contain genuine discussions between developers, which may represent concerns about CI practices if they are pertinent to the project.
We extracted relevant text excerpts from the PR comments related to CI monitoring practices. We then conducted open coding by thoroughly examining the comments and identifying 417 codes. The coded data was analyzed to draw conclusions and insights regarding CI monitoring practices and the need for such monitoring. The Open Coding and Thematic Analysis was performed manually by the first author, without the support of automated tools, to ensure a comprehensive understanding of the data. The core steps of the process involved (1) initial open coding, where the comments were read in detail and codes were generated to represent key ideas; (2) searching for patterns, relationships, and similarities among the codes; and (3) the development of themes by grouping codes based on conceptual similarities. We identified nine key themes during our thematic analysis: (1) build duration, (2) code quality, (3) coverage, (4) hardware, (5) lead time for changes, (6) security, (7) build health and time to fix a broken build, (8) change failure rate and (9) deploy frequency. The themes were chosen to represent the overarching concepts that connect the codes within each theme. We group themes into 2 high-level themes: (i) “Monitoring Opportunities”, which contains mentions of CI metrics but with no indication of monitoring those metrics (e.g. there are recurrent build failures, but no indication as to whether developers know the current or tolerated level of build failures) and (ii) “Monitoring Achievements”, which contains evidence that some CI metrics are monitored (e.g., automatic comments generated by the coverage analysis tools).
@@
The extracted codes and themes were initially assigned by the first author. To reduce the probability of errors inherent in this type of subjective analysis, the second author independently applied the coding guide to label all examples. Finally, we calculated the agreement score between the two sets of labels using the Cohen Kappa statistical coefficient cohen1968weighted , which measures the agreement between two “judges”. We obtained a score of 0.68 (94.94% of agreement), indicating “Substantial Agreement”.
@@
Motivation: In light of the observed level of concern among developers regarding monitoring CI practices in RQ1, we delve deeper into this inquiry in RQ2 to directly solicit developers’ opinions on the importance of monitoring CI practices by conducting a survey. We aim to address questions such as: Do developers prioritize monitoring their CI practices? Do they adhere to specific monitoring processes? Through RQ2, we seek to understand developers’ perceptions of CI practice monitoring and check whether they recognize its importance. Additionally, encourage discussion, knowledge sharing, and community building around best CI practices.
@@
In Step 8 of Figure 1, to identify bug-related issues, we first collected labels associated with project issues. This approach is similar to Vasilescu et al. VasilescuQualityProdOutcomes2015 , where we manually check our projects to understand how they use labels to indicate the presence of bugs. We download all the labels from the 121 projects and checked their description to understand whether the label represented a bug. The resulting list of labels representing bug-related issues is: “defect”, “error”, “bug”, “issue”, “mistake”, “incorrect”, “fault”, “flaw”, “crash”, “bugfix” and “regression.” Similar to the work of Vasilescu et al., VasilescuQualityProdOutcomes2015 , we lowercased the labels and apply the Porter stemming algorithm  [^7] https://tartarus.org/martin/PorterStemmer/ to compare the bug-related labels. Finally, bug-related issues were analyzed to determine correlations between CI practices and the number of bugs in the projects. This data was used in an exploratory survey question to ascertain whether developers observed any correlations between CI practices and the incidence of bugs in the project.